# Language-Guided World Models [Uncaptioned image] A Model-Based Approach to AI Control

### v1

@@ Abstract
Installing probabilistic world models into artificial agents opens an efficient channel for humans to communicate with and control these agents. In addition to updating agent policies, humans can modify their internal world models in order to influence their decisions. The challenge, however, is that currently existing world models are difficult for humans to adapt because they lack a natural communication interface. Aimed at addressing this shortcoming, we develop Language-Guided World Models (LWMs), which can capture environment dynamics by reading language descriptions. These models enhance agent communication efficiency, allowing humans to simultaneously alter their behavior on multiple tasks with concise language feedback. They also enable agents to self-learn from texts originally written to instruct humans. To facilitate the development of LWMs, we design a challenging benchmark based on the game of Messenger (Hanjie et al., 2021), requiring compositional generalization to new language descriptions and environment dynamics. Our experiments reveal that the current state-of-the-art Transformer architecture performs poorly on this benchmark, motivating us to design a more robust architecture. To showcase the practicality of our proposed LWMs, we simulate a scenario where these models augment the interpretability and safety of an agent by enabling it to generate and discuss plans with a human before execution. By effectively incorporating language feedback on the plan, the models boost the agent performance in the real environment by up to three times without collecting any interactive experiences in this environment.
Project website: language-guided-world-model.github.io
Figure 1: Language-guided world models can significantly improve the interpretability, safety, and helpfulness of artificial agents. (a) These models enable an agent to compose intuitive plans and invite a human supervisor to validate and revise those plans. Moreover, they offer the supervisor various strategies for revising a plan: they can provide action-correcting feedback to update the agent’s policy, or language feedback that describes the environment to modify its world model. (b) We design an effective architecture for language-guided world models which exhibits strong compositional generalizability. Our approach converts a trajectory into a long sequence of tokens and trains a Transformer to auto-regressively generate these tokens. It implements a specialized attention mechanism inspired by Hanjie et al. (2021) to incorporate textual information into the observation tokens.
@@ 1 Introduction
Model-based agents are artificial agents equipped with probabilistic “world models” that are capable of foreseeing the future state of an environment (Deisenroth & Rasmussen, 2011; Schmidhuber, 2015). This endows the agents with the ability to plan and learn in imagination (i.e., internal simulation) and has led to exciting results in the field of reinforcement learning (Finn & Levine, 2017; Ha & Schmidhuber, 2018; Chua et al., 2018; Hafner et al., 2023). Most previous work has leveraged this approach to improve the autonomous performance of artificial agents.
In this paper, we endorse and enhance the model-based approach for a different purpose—strengthening the controllability of artificial agents. A model-based agent comprises a policy and a world model, where the policy is optimized with respect to the world model. This modular design can result in more effective control over “model-free” counterparts, allowing humans to alter the agent’s behavior both explicitly, by directly updating its policies (e.g., through reinforcement or imitation learning), or implicitly, by adapting its world model (which subsequently also changes the policies). The latter mechanism has the potential to influence the agent performance across all tasks in the environment, saving substantial human communication effort. For example, telling a cleaning robot that “the floor is slippery” can effectively remind it to handle every object in a room with greater caution.
Current model-based agents are difficult for humans to adapt because they employ purely observational world models. These models can be modified only through observational data, which is not a suitable medium for humans to convey complex intentions (Sumers et al., 2023; Zheng et al., 2023).
@@
To overcome the limitations of observational world models, we develop Language-Guided World Models (LWMs)—world models that can be effectively steered through human verbal communication. Agents equipped with LWMs inherit all the benefits of model-based agents while also being able to incorporate human language-based supervision. Language-based learning reduces human effort, as well as the amount of exploratory observations that these agents need to collect in an environment, mitigating the risk of them taking harmful actions due to insufficient knowledge. LWM-based agents can also self-improve by reading “free” texts composed for guiding humans (e.g., game manuals), lowering subsequent effort to fine-tune them through direct interaction.
Building these models poses a unique research challenge of how to accurately ground language to aspects of an environment’s dynamics, which has been understudied in previous work. This problem is tremendously difficult because it requires a model to understand diverse linguistic concepts (e.g., motion, interaction, appearance, spatial/temporal relation, etc.) that can possibly be used to describe the world. To address this problem, we first construct a benchmark based on the Messenger work of Hanjie et al. (2021), which constructs environments whose dynamics are described by language descriptions depicting attributes of environment entities. We design settings of increasing difficulty, requiring a world model to generalize compositionally with respect to both the language input and the environment dynamics. Our learning approach starts by converting a training trajectory into a token-based representation that is convenient for the integration of the dynamics features extracted from the descriptions. We find that the prominent Transformer model architecture (Vaswani et al., 2017) is ineffective in learning to translate language descriptions into such a representation. Its attention mechanism lacks inductive biases to learn the right feature-extraction function and instead overfits to spurious correlations. We implement a new attention mechanism inspired by the Entity Mapper with Multi-Modal Attention (Hanjie et al., 2021), which mimics a two-step reasoning process. Our results confirm the effectiveness of this architecture, as it robustly generalizes even on the hardest evaluation setting, outperforming various Transformer-based baselines by substantial margins. It is even competitive with an oracle model with a perfect semantic-parsing capability.
@@
Besides conducting intrinsic evaluation, we illustrate a promising application of LWMs. Concretely, we simulate a cautious agent that, instead of performing a task right away, uses its LWM to generate an execution plan and asks a human to review it. This form of pre-execution communication can potentially improve the agent’s interpretability and safety. Moreover, the performance of the agent can also be enhanced by incorporating human feedback on the plan. We train all model-based agents with action-correcting feedback via imitation learning. Our LWM-based agent has the advantage of also being able to incorporate language feedback describing the environment dynamics. We demonstrate that the language understanding capabilities of our LWM are sufficient to improve the agent’s capabilities dramatically. In our hardest evaluation setting, without gathering additional interactions in the real environment, this agent achieves an average cumulative reward that is up to three times that of an agent trained without language supervision.
@@
We hope that our work will serve as a catalyst for exploring novel approaches that enable humans to tap into the world models of artificial agents. More generally, we call for the design of modular agents whose components are parameterized by natural language. We hypothesize that such agents would allow for effective incorporation of diverse linguistic signals from humans, potentially surpassing the currently prevalent model-free approaches that use language that aims at only controlling a policy (i.e. instructions) (Bisk et al., 2016; Misra et al., 2018; Anderson et al., 2018) or feed language conveying diverse intentions directly to a black-box policy (e.g., Narasimhan et al. (2018); Hanjie et al. (2021); Zhong et al. (2021) and work on large language models like Ouyang et al. (2022)).
@@ 2 Background: world models
We consider a Markov Decision Process (MDP) environment $E$ with state space ${\mathcal{S}}$, action space ${\mathcal{A}}$, and transition function $M(s_{t+1},r_{t+1},d_{t+1}\mid s_{t},a_{t}):{\mathcal{S}}\times{\mathcal{A}}\rightarrow\Delta({\mathcal{S}}\times\mathbb{R}\times\{0,1\})$, where $\Delta$ denotes the set of all probability distributions over a set. An agent implementing a policy $\pi(a\mid s):{\mathcal{S}}\rightarrow\Delta({\mathcal{A}})$ interacts with the environment by choosing actions using its policy. Taking an action $a_{t}\sim\pi(s_{t})$ in state $s_{t}$ transitions the agent to a new state $s_{t+1}$, and incurs a reward $r_{t+1}$ and a termination signal $d_{t+1}$, where $s_{t+1},r_{t+1},d_{t+1}\sim M(s_{t},a_{t})$.
A (one-step) world model (Robine et al., 2023; Micheli et al., 2023; Hafner et al., 2023) approximates $M(s_{t+1},r_{t+1},d_{t+1}\mid s_{t},a_{t})$. A model-based agent learns this model in addition to its policy and leverages it to plan or optimize the policy. To improve this type of agent, being able to adapt only the policy is insufficient. Learning an optimal policy with respect to an erroneous world model still results in suboptimal behavior in the real world. Therefore, it is essential to be able to also alter the world model.
@@
Figure 2: Messenger environment.
Relying on observation-based adaptation leads to two major drawbacks of observational world models. First, humans cannot effectively control them, because observations are inadequate for conveying complex, abstract human intentions. Second, collecting observations requires taking real actions in the environment, which can be expensive, time-consuming, and risky. We elucidate these drawbacks using the game of Messenger (Hanjie et al., 2021). In an example game shown in Figure 2, the ferry plays the role of an enemy who will immediately purge the player and end the game once it collides with them. An observational world model cannot determine the role of this entity without witnessing it killing the player, which is an undesirable event to trigger. Past experiences could be misleading: for example, the model might have seen the ferry playing only other roles in the past, and therefore infer that it is unlikely to be an enemy. The game manual reveals the identity of the enemy, but the model is unable to utilize such information. The only way for a human to communicate the enemy role to this model is to draw observations illustrating the player’s death after colliding with the entity, which requires a lot of manual effort.
@@ 3 Language-guided world models (LWMs)
LWMs can interpret language descriptions to adapt to new environment dynamics. This class of models addresses the two aforementioned drawbacks of observational world models. They allow humans to easily adapt their behavior through natural language. Consequently, humans can effectively assist these models by providing language-based information, significantly reducing the amount of interactive experiences needed in real environments. They can also leverage texts written for humans to quickly learn new environment dynamics.
@@ 3.1 Formulation
Our formal setup augments the MDP setting by assuming that the environment is accompanied by a language manual =$\ell=\{l_{1},\cdots,l_{N}\}$ consisting of text descriptions $l_{i}$ that disclose information about the transition function. Our goal is to learn a model $M_{\theta}(s_{t+1},r_{t+1},d_{t+1}\mid h_{t},\ell)$ that can interpret the manual $\ell$ to better predict $s_{t+1}$, $r_{t+1}$, and $d_{t+1}$. We first present the general training and evaluation procedures in this section and describe the specific architecture that we implemented in the next section.
Training.
@@
Let $P_{{\mathcal{S}},{\mathcal{A}}}^{\textrm{train}}(E)$ be a distribution over manual-augmented environments with state space ${\mathcal{S}}$ and action space ${\mathcal{A}}$. The training data for our world model is a set of tuples ${\mathcal{B}}_{\textrm{train}}=\{(\tau^{i},\ell^{i})\}_{i=1}^{|{\mathcal{B}}_{\textrm{train}}|}$ where $\tau^{i}$ is a trajectory generated in an environment $E^{i}\sim P_{{\mathcal{S}},{\mathcal{A}}}^{\textrm{train}}$ and $\ell^{i}$ is the manual accompanying that environment. Each trajectory $\tau=(s_{1},r_{1},d_{1},a_{1},\ldots,s_{T},r_{T},d_{T})$ is a sequence of states, actions, rewards, and termination signals. These trajectories are generated using a behavior policy $\pi$, which can be a rule-based or learned policy, or a human. We train an LWM as an auto-regressive sequence generator to minimize the cross-entropy loss on the training data
$\min_{\theta}J(\theta;{\mathcal{B}}_{\textrm{train}})\triangleq\min_{\theta}-\frac{1}{|{\mathcal{B}}_{\textrm{train}}|}\sum_{(\tau,\ell)\in{\mathcal{B}}_{\textrm{train}}}\frac{1}{T_{\tau}-1}\sum_{t=1}^{T_{\tau}-1}\log M_{\theta}(s^{\tau}_{t+1},r^{\tau}_{t+1},d^{\tau}_{t+1}\mid h^{\tau}_{t},\ell)$ (1)
where $T_{\tau}$ is number of time steps in $\tau$ and $h^{\tau}_{t}$ is a (learned) representation of $\tau$ up to time step $t$.
Evaluation.
Previous work mostly evaluates world models on downstream control tasks (Ha & Schmidhuber, 2018; Finn & Levine, 2017; Hafner et al., 2023). While it serves a practical end goal, this extrinsic evaluation scheme does not truly measure the quality of a world model since the result also depends on the strength of the search or RL algorithm used. Since our focus is to learn accurate world models that can be employed for different purposes (including control tasks), we also conduct intrinsic evaluation of the models. We estimate the probabilistic divergence with respect to the true model by computing the approximate cross entropy $J(\theta;{\mathcal{B}}_{\textrm{eval}})$, where ${\mathcal{B}}_{\textrm{eval}}$ contains trajectories generated in previously unseen environments. [^1] See Appendix E for a precise mathematical interpretation of this metric. We also use the learned model to generate imaginary trajectories and compute the precision of the next state, reward, and termination-signal predictions.
3.2 Model
Learning LWMs poses a challenging problem that involves the retrieval and incorporation of information expressed in different modalities. Due to the complexity of the general problem, we consider a simplified setting in which the manual conveys information about attributes of entities in an environment, and the environment state follows an entity-based representation. Despite this simplification, our setting remains highly general and is by no means easy to tackle. As we will show, the state-of-the-art Transformer architecture struggles to generalize compositionally in this setting. In this section, we present a more effective modeling approach. Our proposed architecture is illustrated in 0(b).
Entity-based state representation.
We focus on 2D grid-world environments featuring a set of $C$ entities. Each entity $c$ has an identity (e.g., a mage) and $K$ attributes (e.g., movement pattern, role). Each identity is represented by an index $i\in[I]$, where $I$ is the number of identities. Similarly, each attribute value is an index $v_{k}\in[V_{k}]$ where $V_{k}$ is the number of distinct values of attribute $k$. The identity and attributes of an entity vary among environment instances. The entities can travel in an environment and interact with one another, and the outcomes of their interactions are decided on the basis of their attributes.
A state $s$ in these worlds is represented by an $H\times W$ grid with $C$ channels (an $H\times W\times C$ tensor), where each channel corresponds to a single entity. In each channel $c$, there is a single non-zero cell $s(h,w,c)\in[I]$ that represents the identity and position of entity $c$.
World modeling as sequence generation.
Our model is an encoder-decoder Transformer which encodes a manual $\ell$ and decodes a trajectory $\tau$. We transform the trajectory into a long sequence of tokens and train the model as a sequence generator. Prior work (Robine et al., 2023; Micheli et al., 2023; Hafner et al., 2023) employs a latent-variable model to discretize observations into tokens. However, for this problem, to effectively incorporate information extracted from the manual, we require an observation representation that disentangles the entities (i.e., the states and attributes of the entities must be captured by disjoint sets of tokens). Such a representation is easily derived for entity-based observations without having to implement a latent-variable model. We defer the problem of learning entity-disentangled representation for pixel-based observations to future work.
Our model processes a data point $(\tau,\ell)$ as follows. For the manual =$\ell=\{l_{i}\}_{i=1}^{N}$, we first use a pre-trained BERT model to convert each description $l_{i}$ into a sequence of hidden vectors. We feed each sequence to a Transformer encoder, which outputs a tensor $\bm{m}^{\textrm{enc}}$ of size $N\times L\times D$, where $N$ is the number of descriptions, $L$ is the maximum number of words in a description, and $D$ is the hidden size.
For the trajectory, we convert each tuple $(a_{t-1},s_{t},r_{t},d_{t})$ into a token block $B_{t}$. The first action $a_{0}$ is set to be a special <s> token. Each state $s_{t}$ is mapped to $3C$ tokens $(i_{t}^{1},h^{1}_{t},w^{1}_{t},\cdots,i^{C}_{t},h^{C}_{t},w^{C}_{t})$, which represents each of the $C$ entities by its identity $i$ followed by its coordinates $(h,w)$ in the grid. The real-valued reward $r_{t}$ is discretized into an integer label, and the termination signal $d_{t}$ is translated into a binary label. In the end, $B_{t}$ consists of $33C+$ 3 tokens $(a_{t-1},i_{t}^{1},h^{1}_{t},w^{1}_{t},\cdots,i^{C}_{t},h^{C}_{t},w^{C}_{t},r_{t},d_{t})$. Finally, we concatenate all $T_{\tau}$ blocks in the trajectory into a sequence of $T_{\tau}\times(3C+3)$ tokens, embed them into a $T_{\tau}\times(3C+3)\times D$ tensor, and add positional embeddings. We will use bold notation (e.g. $\bm{a}$, $\bm{i}$) to refer to the resultant embeddings of the tokens.
Entity mapper with multi-modal attention.
We implement a variant of EMMA (Hanjie et al. (2021)) that first identifies the manual description that mentions each entity and extracts from it words corresponding to the attributes of the entity. Formally, from the tensor $\bm{m}^{\textrm{enc}}$ computed by the encoder, we generate a key tensor $\bm{m}^{\textrm{key}}$ and a value tensor $\bm{m}^{\textrm{val}}$, both of which are of size $N\times L\times D$, where
$\bm{m}^{\textrm{key}}_{n}$ $=\texttt{Softmax}(\texttt{Linear}_{\textrm{key}}^{D\rightarrow 1}(\bm{m}_{n}^{\textrm{enc}})^{\top})\bm{m}_{n}^{\textrm{enc}}$ (2)
$\bm{m}^{\textrm{val}}_{n}$ $=\texttt{Softmax}(\texttt{Linear}_{\textrm{val}}^{D\rightarrow 1}(\bm{m}_{n}^{\textrm{enc}})^{\top})\bm{m}_{n}^{\textrm{enc}}$ (3)
for $1\leq n\leq N$. Here, $\texttt{Linear}^{D\rightarrow 1}_{\textrm{key}}$ and $\texttt{Linear}^{D\rightarrow 1}_{\textrm{val}}$ are linear layers that transform the input’s last dimension from $D$ to 1, and $\texttt{Softmax}(\cdot)$ applies the softmax function to the last dimension. Intuitively, we want each $\bm{m}^{\textrm{key}}_{n}$ to retain words that signal the identity of the entity mentioned in the $n$-th description (e.g., ferry, plane, researcher), and $\bm{m}^{\textrm{val}}_{n}$ to capture words that convey the attributes of the entity (e.g., approaching, deadly, fleeing).
Let $\bm{i}^{c}_{t}$ be the embedding of the identity of entity $c$. We perform a dot-product attention with $\bm{i}^{c}_{t}$ as the query, $\bm{m}^{\textrm{key}}$ as the set of keys, and $\bm{m}^{\textrm{val}}$ as the set of values to compute the attribute features $\bm{z}^{c}_{t}$ of entity $c$
The features are incorporated to the input of the decoder Transformer at each time step $t$ as follows:
$(\bm{a}_{t-1},{\color[rgb]{1,0,0}\bm{i}_{t}^{1}+\bm{z}^{1}_{t}},\bm{h}^{1}_{t},\bm{w}^{1}_{t},\cdots,{\color[rgb]{1,0,0}\bm{i}^{C}_{t}+\bm{z}^{C}_{t}},\bm{h}^{C}_{t},\bm{w}^{C}_{t},\bm{r}_{t},\bm{d}_{t})$ (5)
Unlike a standard encoder-decoder Transformer model, our model does not perform cross-attention between the encoder and the decoder because information from the encoder has already been incorporated into the decoder through EMMA.
ChatGPT-based hard attention.
We also attempt an alternative approach that leverages the language-understanding capabilities of ChatGPT. Through few-shot prompting, we instruct this model to determine the identity of the entity mentioned in each manual description. In this approach, we generate only the set of values $\bm{m}^{\textrm{val}}$ as in Eq 3. Instead of learning soft attention, we directly route the values to the identity embeddings. Concretely, the feature vector added to $\bm{i}^{c}_{t}$ in Eq 5 is $\bm{z}^{c}_{t}=\bm{m}^{\textrm{val}}_{\hat{i}_{c}}$ where $\hat{i}_{c}\in[N]$ is the index of the description that mentions entity $c$ according to ChatGPT.
Model training.
We train the model to minimize cross-entropy loss with respect to ground-truth (tokenized) trajectories in the training set. The label at each output position is the next token in the ground-truth sequence. Particularly, we do not compute the losses at the positions of the action tokens and the first block’s tokens, because those tokens will be set during inference.
4 The Messenger-Wm benchmark: Evaluating the compositional generalizability of world models
We suppose that the dynamics of an environment is governed by independent latent factors. A robust world model must adapt robustly to changes in the values of these factors, generalizing to novel compositions of these values. We construct a benchmark to evaluate this capability of world models.
Environment.

### v2

@@ Abstract
This paper introduces the concept of Language-Guided World Models (LWMs)—probabilistic models that can simulate environments by reading texts. Agents equipped with these models provide humans with more extensive and efficient control, allowing them to simultaneously alter agent behaviors in multiple tasks via natural verbal communication. In this work, we take initial steps in developing robust LWMs that can generalize to compositionally novel language descriptions. We design a challenging world modeling benchmark based on the game of Messenger (Hanjie et al., 2021), featuring evaluation settings that require varying degrees of compositional generalization. Our experiments reveal the lack of generalizability of the state-of-the-art Transformer model, as it offers marginal improvements in simulation quality over a no-text baseline. We devise a more robust model by fusing the Transformer with the EMMA attention mechanism (Hanjie et al., 2021). Our model substantially outperforms the Transformer and approaches the performance of a model with an oracle semantic parsing and grounding capability. To demonstrate the practicality of this model in improving AI safety and transparency, we simulate a scenario in which the model enables an agent to present plans to a human before execution, and to revise plans based on their language feedback.
Language-Guided World Models [Uncaptioned image]
A Model-Based Approach to AI Control
thanks: First two authors contribute equally. Correspondence email: kxnguyen@berkeley.edu.Alex Zhang, 11footnotemark: 1Khanh Nguyen, Jens Tuyls, Albert Lin, Karthik Narasimhan Princeton University  University of California, Berkeley University of Southern California Project website: language-guided-world-model.github.io
Figure 1: Language-guided world models (LWMs) offer human an efficient mechanism to regulate artificial agents. (a) We illustrate a potential application of LWMs to improving AI safety and transparency. These models enable an agent to generate visual plans and invite a human supervisor to validate them. Moreover, the human can adjust the plans by modifying the agent’s world model with language feedback, in addition to directly correcting its policy. (b) We design an architecture for LWMs that exhibits strong compositional generalization. We replace the cross-attention mechanism of the standard Transformer with a new attention mechanism inspired by Hanjie et al. (2021) to effectively incorporate language descriptions. We then train a model that auto-regressively generates tokenized observations conditioned on language descriptions and actions.
@@ 1 Introduction
Model-based agents are artificial agents equipped with probabilistic “world models” that are capable of foreseeing the future state of an environment (Deisenroth and Rasmussen, 2011; Schmidhuber, 2015). World models endow these agents with the ability to plan and learn in imagination (i.e., internal simulation) and have led to exciting results in the field of reinforcement learning (Finn and Levine, 2017; Ha and Schmidhuber, 2018; Chua et al., 2018; Hafner et al., 2023). These models have been studied extensively for the purpose of improving the autonomous performance of artificial agents.
In this paper, we endorse and enhance the model-based approach for a different goal: to strengthen the controllability of artificial agents. Since all policies of a model-based agent are optimized with respect to a common world model, a human can adjust multiple policies simultaneously by making appropriate changes to this model. This mechanism complements the model-free approach that updates policies individually, offering greater efficiency and flexibility in control. For example, by incorporating the fact that the floor is slippery into the world model of a robot, a person can effectively remind it to handle every object in a room with greater caution. If the performance of the robot on a task remains unsatisfactory, the person can continue to fine-tune its policy for that specific task. In contrast, without a world model, they have to separately adapt the robot’s policies to the slippery-floor condition.
@@
The model-based approach requires world models that can be easily modulated by humans. Traditional world models fall short in this quality because they can only be modified using observational data, which is not a suitable medium for humans to convey intentions (Sumers et al., 2023; Zheng et al., 2023). To overcome the limitations of these models, we develop Language-Guided World Models (LWMs)—world models that can be effectively steered through human verbal communication. Agents equipped with LWMs inherit all the benefits of model-based agents while being able to incorporate language-based supervision. This capability reduces human teaching effort and mitigates the risk of agents taking harmful actions in an environment to explore its dynamics. LWM-based agents can also self-improve by reading “free” texts composed to guide humans (e.g., game manuals), reducing the subsequent effort to fine-tune them through direct interaction.
Building LWMs poses a unique research challenge: grounding language to environmental dynamics. This problem is difficult because the language used to describe environment dynamics can be incredibly rich and complex, encompassing a wide range of concepts such as entity names, appearances, motions, interactions, spatial and temporal relations, and more. Moreover, in natural settings, especially when describing artificial environments (e.g., games), new concepts are often introduced but may not always be clearly defined. Humans deal effectively with this issue because they possess remarkable reasoning capabilities that allow them to infer word meanings from observations. For example, a caption like “the Ziff, which is chasing the player, is extremely hostile” and a video depicting this scene likely provide enough clues for a person to determine what “the Ziff” refers to, assuming that they are familiar with the concept of “chasing”. Not only understanding word meanings, humans are also capable of applying newly learned words in novel ways, enabling imagination of new dynamics, such as envisioning a “fleeing Ziff” that runs away from the player.
Toward building world models with similar capabilities, we construct a benchmark based on the game of Messenger (Hanjie et al., 2021). In this benchmark, a model is given trajectory “videos” of games involving several entities interacting with each other. Each video is accompanied by language descriptions of the attributes of the entities. The model begins with almost zero language understanding and has to identify the entities and learn the grounded meanings of their attributes purely by watching the videos. At test time, it must demonstrate compositional generalization by being able to simulate environments featuring entities with attributes different from those it observes during training. For example, it has to portray a “fleeing mage” despite having only seen the mage chase the player in training games. We design three evaluation settings that test for incrementally greater degree of compositional generalization.
Despite its apparent simplicity, our benchmark covers many complications in building robust LWMs. We find that the prominent Transformer model (Vaswani et al., 2017) struggles in the harder evaluation settings. Even with a ground-truth disentangled representation of the observations, the model cannot learn generalizable grounding functions and yields minimal improvements in simulation quality compared to a model that ignores the language descriptions entirely. We augment the model with the EMMA attention (Hanjie et al., 2021), which mimics a two-step reasoning process. Our results confirm the effectiveness of this new architecture, as it robustly generalizes even in the hardest evaluation setting, outperforming baselines by substantial margins in various evaluation metrics. It is even competitive with a skyline model with an oracle semantic parsing and grounding capability.
@@
Last but not least, we illustrate a promising application of LWMs by simulating a cautious agent that, instead of performing a task right away, uses its LWM to generate an execution plan and asks a human to review it (1(a)). This form of pre-execution communication can potentially improve the agent’s safety and transparency, following the spirit of the guaranteed safe AI approach proposed by Dalrymple et al. (2024). Moreover, it allows the human to improve the performance of the agent by revising the plan. In this setting, our LWM-based agent has the advantage of being able to assimilate language feedback describing the environment dynamics. We demonstrate that the language understanding capabilities of our proposed LWM are sufficient to enact this strategy. In the most challenging evaluation setting, without gathering additional interactions in the environment, the agent equipped with our model achieves an average reward three to four times higher than that of an agent using an observational world model.
@@
We hope that our work will serve as a catalyst for exploring novel approaches to developing robust language-guided world models. More generally, we call for the design of modular agents whose components are parameterized by natural language. As previously argued, a modular design can dramatically boost communication efficiency, because the same component may be involved in the learning of various policies. We hypothesize that this approach can potentially surpass the efficiency of the currently prevalent approach that integrates language into a monolithic policy (e.g, Bisk et al. (2016); Misra et al. (2018); Anderson et al. (2018); Narasimhan et al. (2018); Hanjie et al. (2021); Zhong et al. (2021) and work on large language models like Ouyang et al. (2022)).
@@ 2 Background: world models
We consider a Markov Decision Process (MDP) environment $E$ with state space ${\mathcal{S}}$, action space ${\mathcal{A}}$, and transition function $M:{\mathcal{S}}\times{\mathcal{A}}\rightarrow\Delta({\mathcal{S}}\times\mathbb{R}\times\{0,1\})$, where $\Delta$ denotes the set of all probability distributions over a set. An agent implementing a policy $\pi(a\mid s):{\mathcal{S}}\rightarrow\Delta({\mathcal{A}})$ interacts with the environment by choosing actions using its policy. Taking an action $a_{t}\sim\pi(s_{t})$ in state $s_{t}$ transitions the agent to a new state $s_{t+1}$, and incurs a reward $r_{t+1}$ and a termination signal $d_{t+1}$, where $s_{t+1},r_{t+1},d_{t+1}\sim M(s_{t},a_{t})$.
A (one-step) world model $M_{\theta}$ (Robine et al., 2023; Micheli et al., 2023; Hafner et al., 2023) is an approximation of $M(s_{t+1},r_{t+1},d_{t+1}\mid s_{t},a_{t})$. A model-based agent uses data gathered in the environment to construct a world model and leverages it to learn policies for accomplishing tasks. [^1] Note that $M_{\theta}$ includes a reward function but can be combined with any other reward function for learning. In contrast, a model-free agent learns its policies directly from data collected in the environment.
Model-based agents can require less effort to adapt.
Because all policies of a model-based agent are derived from a shared world model, any modifications made to this model would affect all of them. This feature can be exploited to reduce human effort in controlling this type of agent. Specifically, suppose we concern $m$ tasks in the environment, necessitating $m$ policies. If there is a change in the environment dynamics, a model-based agent only needs task-agnostic data to replicate this change in its world model. It can then re-optimize its policies with respect to the updated model. Meanwhile, a model-free agent needs to collect task-specific data to re-train all of its $m$ policies. The data collection cost of the model-free approach scales with $m$, whereas that of the model-based approach is independent of $m$, since the policy re-optimization step uses only data generated by the world model.
Observational world models.
@@
Relying on observation-based adaptation leads to two drawbacks. First, controlling these models is difficult because observations are inadequate for conveying complex, abstract human intentions. Second, collecting observations requires taking real actions in the environment, which can be expensive, time-consuming, and risky.
@@ 3 Language-guided world models (LWMs)
We introduce LWMs, a new class of world models that can interpret language descriptions to simulate environment dynamics. These models address the drawbacks of observational world models. They allow humans to easily adapt their behavior through natural means of communication. Consequently, humans can effectively assist these models, significantly reducing the amount of interactive experiences that they need to collect in environments. In addition, these models can also leverage pre-existing texts written for humans, saving human effort to fine-tune them.
@@ 3.1 Formulation
We consider a family of environments $E(\bm{v})$ whose transition function has the form $M(s_{t+1},r_{t+1},d_{t+1}\mid h_{t},\bm{v})$ where $\bm{v}$ is a parameter vector. Plugging in a specific $\bm{v}$ gives rise to an environment. We assume that each environment $E(\bm{v})$ is accompanied by a language manual $\bm{\ell}=(l_{1},\cdots,l_{N})$ consisting of language descriptions $l_{i}$. This manual describes $\bm{v}$ and the internal operations of $M$. Our goal is to learn a world model $M_{\theta}(s_{t+1},r_{t+1},d_{t+1}\mid h_{t},{\color[rgb]{.75,0,.25}\bm{\ell}})$ that approximates the true dynamics $M(s_{t+1},r_{t+1},d_{t+1}\mid h_{t},\bm{v})$.
@@
The training data for our LWMs is a dataset $\{(\tau^{i},\bm{\ell}^{i})\}$ where $\tau^{i}$ is a trajectory generated in an environment $E(\bm{v}_{i})$ with $\bm{v}_{i}$ drawn from some distribution $P_{\textrm{train}}$, and $\bm{\ell}^{i}$ is the accompanying manual. Each trajectory $\tau=(s_{1},r_{1},d_{1},a_{1},\ldots,s_{T},r_{T},d_{T})$ is a sequence of states, actions, rewards, and termination signals. It can be viewed as a “video” that is annotated with actions and rewards. The trajectories are generated using a behavior policy, which can be a rule-based or learned policy, or a human.
3.2 Modeling entity-based environments
We view an environment as a set of $C$ entities interacting with each other within a constrained space. Each entity $c$ has a set of $K$ attributes, each of which has value $v_{k}^{c}$. There is a special attribute called the identity of the entity (e.g., the name of a character or object in a video game). Each action triggers an event that changes a subset of attributes of a group of entities. The specific change is determined by the attributes of the entities involved in the event (e.g., an enemy entity attacks a player when colliding with them). In this work, we assume that each description in a manual portrays all attributes of an entity; hence, the number of descriptions $N$ is equal to $C$.
Testing for compositional generalization.
[…]
This problem requires a LWM to be able to learn a representation of the transition function $M(\bm{v})$ by studying the language of the manuals, and to extract the specific parameters $\bm{v}$ described by each manual. The function $M(\bm{v})$ has two important properties. The first is the independence among its parameters because they represent orthogonal attributes. The second is the locality of the parameters, as each is an attribute associated with only a single entity. These properties make it difficult to recover the function exactly from purely observational data without injecting strong inductive biases into the learning model.
3.3 The Messenger-WM benchmark
Figure 2: Messenger environment with manual.
The game of Messenger, developed by (Hanjie et al. (2021); Figure 2) exemplifies the class of environments discussed in the previous section. Despite being a simple grid-world environment, the dynamics possess the independence and locality properties that we want to study. In fact, it is our intention to use this visually simplistic environment to highlight the challenges in building LWMs that are orthogonal to the computer graphics challenge of mapping state representations to realistic-looking outputs.
Environment dynamics.